OpenStudy (anonymous):

2x-9+7x-27=90 Can someone please help explain how to work out this problem step-by-step, solving for x?

OpenStudy (jonnyvonny):

Mmk, so, to solve for x, we need to get x alone; by itself. Are you familiar with reverse pemdas?

OpenStudy (anonymous):

First time I have ever heard of it actually, does it have to doing with everything in reverse? Such as addition and subtraction from right to left?

OpenStudy (jonnyvonny):

Exactly

OpenStudy (jonnyvonny):

What math are you in, out of curiosity?

OpenStudy (anonymous):

Geometry, this is part of a larger question to do with angles, but its the equation itself I am confused with.

OpenStudy (jonnyvonny):

We use reverse pemdas to change an equation so we may get what we want; make it easy. Oh, mmk.

OpenStudy (jonnyvonny):

Mmk, so, first off we may simplify that problem. Do you see how?

OpenStudy (anonymous):

2x-9+7x-27=90 So do we take the -27 and subtract it from both sides of the equation? i.e. 2x-9+7x-27=90 +27 -27

OpenStudy (jonnyvonny):

Exactly, you get it. That is perfectly great. Personally, I would subtract the -9 and -27 first, then do it, but it is up to you. We will have to do that eventually, though.

OpenStudy (jonnyvonny):

And add the 2x and 7x.

OpenStudy (jonnyvonny):

But yeah, you got it man. Now, solve it while I do it aswell, and see what we get; see if it is the same.

OpenStudy (anonymous):

Alrighty

OpenStudy (jonnyvonny):

What did you get?

OpenStudy (jonnyvonny):

If you need some more help, feel free to say so.

OpenStudy (anonymous):

Just a sec, I'm writing it out by hand. Showing work on this assignment is key, apparently. as with all math.

OpenStudy (jonnyvonny):

No problem.

OpenStudy (anonymous):

I got 9.1 Something tells me that's not right.

OpenStudy (jonnyvonny):

I got 14. So, lets work this out.

OpenStudy (jonnyvonny):

2x-9+7x-27=90

OpenStudy (anonymous):

Alright, so i ended up starting with taking the 27 off both sides of the equation like before, but what would you recommend starting with?

OpenStudy (jonnyvonny):

Personally, combining like terms

OpenStudy (jonnyvonny):

Combining the -27 and -9 and 2x and 7x.

OpenStudy (jonnyvonny):

It will result in this: 9x-36=90

OpenStudy (jonnyvonny):

From then, what do?

OpenStudy (jonnyvonny):

What do you think we should do, is what I mean to ask.

OpenStudy (anonymous):

Work the 36 out of the equation, correct?

OpenStudy (jonnyvonny):

Correct. And what new equation will we get once you do so?

OpenStudy (anonymous):

If I did everything properly, 9x=126.

OpenStudy (jonnyvonny):

Yup, now divide the 9 over, and we should get the right answer

OpenStudy (anonymous):

9/26= 14 Well dang on. Awesome.

OpenStudy (jonnyvonny):

So, when you did it your way, I'm assuming that after you added +27 on both sides, you accidentally added a -9 to both sides, instead of a +9.

OpenStudy (anonymous):

Haha. Yep. That's exactly what I had done.

OpenStudy (jonnyvonny):

Yeah, it still happens to me lol. When your answer blows up on you, either you don't know what you're doing at all, or there was a minor error. If you have any more questions, feel free to ask.

OpenStudy (anonymous):

Alright, thanks. I have one more like question, and if that does the same, I'll ask. Thanks for the help man.

OpenStudy (jonnyvonny):

No problem. Kill that question man; you got it.